# REOrdering Patches Improves Vision Models

### v1

@@ 1 Introduction
However, this assumption does not hold in the context of modern, long-sequence models which introduce strong inductive biases such as locality beltagy2020longformerlongdocumenttransformer , recurrence dai2019transformerxlattentivelanguagemodels , or input-dependent state dynamics gu2024mambalineartimesequencemodeling that are sensitive to input ordering. Through mechanisms such as sparse attention via masking summarizing early parts of a long-sequence into latent representations, these methods are able to model long-sequences in a computationally tractable fashion. However, we assert that these design choices introduce a strong dependency on patch ordering, something that has previously been overlooked.
In this work, we demonstrate that merely swapping the row-major scan for one of six studied alternatives, such as column-major or Hilbert curves, yields measurable accuracy gains across multiple long-sequence backbones. Further, we introduce REOrder, a method to learn an optimal patch ordering. REOrder first approximates which patch ordering may lead to best performance and then learns to rank patches in order of importance with reinforcement learning. Specifically, to convert the 2-D image into a 1-D sequence, we first quantify the compressibility resulting from each of six different patch orderings. Then, a Plackett-Luce ranking model is initialized with the least compressible ordering and further trained to minimize classification loss with REINFORCE. REOrder improves performance on ImageNet-1K by up to $3.01\%~{}(\pm 0.23\%)$ and Functional Map of the World by $13.35\%~{}(\pm 0.21\%)$. Code and animations are available on the project page.
@@ 2 Related Works
As established, the quadratic computational cost of full self-attention with respect to sequence length makes processing long-sequences of patches very expensive for standard Vision Transformers. To address this, significant research has focused on developing efficient transformer architectures and long-sequence models that reduce this complexity, such as Sparse Attention child2019generatinglongsequencessparse , Longformer beltagy2020longformerlongdocumenttransformer , Transformer-XL dai2019transformerxlattentivelanguagemodels , and Mamba gu2024mambalineartimesequencemodeling ; renAutoregressivePretrainingMamba2024 . While making modeling tractable, these methods often introduce specific inductive biases in how they process sequences, potentially making them sensitive to the order of the input tokens (which we discuss further in Section 3.3). In this work, we study the effect of patch ordering on Longformer, Transformer-XL, and Mamba, demonstrating that the inductive biases inherent in these different long-sequence modeling approaches lead to substantial accuracy variance across different patch orders.
@@ 3.1 Self-attention and permutation equivariance
$X=[x_{1};\dots;x_{n}]\in\mathbb{R}^{n\times d}.$
@@
$\operatorname{Attn}(X)=\operatorname{softmax}\!\Bigl{(}\tfrac{(XW_{q})(XW_{k})^{\!\top}}{\sqrt{d}}\Bigr{)}\,XW_{v},\qquad W_{q},W_{k},W_{v}\in\mathbb{R}^{d\times d}$ (1)
where $W_{q}$, $W_{k}$, and $W_{v}$ are learnable query, key, and value matrices, respectively.
@@
For every permutation matrix $P\in\{0,1\}^{n\times n}$,
$\operatorname{Attn}(PX)=P\,\operatorname{Attn}(X).$ (2)
Equation (2) states that self-attention is equivariant to arbitrary permutations of the patch ordering. Hence, the Vision Transformer, composed entirely of self-attention, is also permutation-equivariant. A proof for Proposition (3.1) is in Appendix B; a thorough analysis is conducted by Xu et. al. xuPermutationEquivarianceTransformers2024 .
@@ 3.3 Self-attention approximations and sensitivity to patch order
Recent work reduces the $\mathcal{O}(n^{2})$ cost of full attention by sparsifying the attention pattern, factoring the softmax kernel, or replacing it with a learned recurrence. These changes, however, also break the full permutation-equivariance guarantee of Eq. (2). Below we inspect three representative models. Their attention patterns are visualized in Appendix G.
@@
$Q=\mathbf{H}W_{q},\qquad K=\tilde{\mathbf{H}}\,W_{k},\qquad V=\tilde{\mathbf{H}}\,W_{v}$
Let $P$ permute the $L$ patches of only the current segment (but not the memory). Memory corresponds to previous segments, so its rows are fixed under permutation.
$P\tilde{\mathbf{H}}=[\operatorname{SG}(\mathbf{M});\,P\mathbf{H}]$
Because the second and third logit terms of $M_{\text{TXL}}(i,j)$ depend on the relative index $i-j$, conjugating by $P$ does not commute with the soft-max:
$A^{\text{TXL}}(P\mathbf{H})\neq P\,A^{\text{TXL}}(\mathbf{H}).$
@@
Longformer beltagy2020longformerlongdocumenttransformer uses a sliding-window pattern of width $w$ plus $g$ global tokens with their own projections $W_{q}^{\text{(g)}},W_{k}^{\text{(g)}},W_{v}^{\text{(g)}}$. Because the local mask $M^{\text{local}}$ fixes which $(i,j)$ pairs are valid, applying $P$ to the rows/columns re-labels many entries as illegal ()($\infty$)( - ∞ ) and others as legal, so
$\operatorname{softmax}(PM^{\text{local}}P^{\!\top})\neq P\,\operatorname{softmax}(M^{\text{local}})\,P^{\!\top}.$
The same holds for the global mask unless $P$ preserves the hand-picked global positions. Therefore, Longformer is sensitive to patch ordering by design. It converts $\mathcal{O}(n^{2})$ complexity to $\mathcal{O}(nw)$ under a rigid spatial prior.
@@
$h_{t}\;=\;A_{t}\,h_{t-1}\;+\;B_{t}\,x_{t},\qquad y_{t}=C_{t}\,h_{t},$
with inputs $x_{t}$ produced by scanning the patch sequence left-to-right. A permutation $P$ re-orders the stream, changing $(B_{t},C_{t},\Delta_{t})$ and the sequence of matrix multiplications, so the recurrence yield is different. Therefore Mamba is also permutation sensitive. Its $\mathcal{O}(n)$ complexity comes at the price of a fixed processing order.
In this work, we use ARM as introduced by Ren et. al. renAutoregressivePretrainingMamba2024 . In ARM, every Mamba layer runs four causal scans in different directions $d$ and sums their outputs before the channel-mixing MLP.
For each $d\in\{\!\rightarrow,\leftarrow,\downarrow,\uparrow\}$ let $x_{t}^{(d)}$ be the patch encountered at time step $t$ when traversing the image in direction $d$. The direction-specific recurrence is
$h_{t}^{(d)}\;=\;A_{t}^{(d)}\,h_{t-1}^{(d)}+B_{t}^{(d)}\,x_{t}^{(d)},\qquad y_{t}^{(d)}\;=\;C_{t}^{(d)}\,h_{t}^{(d)},$
@@
$y_{t}\;=\;\sum_{d\in\{\!\rightarrow,\leftarrow,\downarrow,\uparrow\}}y_{t}^{(d)}.$
Similar to Mamba, because each scan is individually directional, an arbitrary permutation $P$ of the patch order both re-orders the input streams $\{x_{t}^{(d)}\}$ and alters the learned parameter sequences $\bigl{\{}A_{t}^{(d)},B_{t}^{(d)},C_{t}^{(d)}\bigr{\}}$. Hence the mapping still violates permutation equivariance $y(PX)\neq P\,y(X)$.
@@ 4 Does Patch Order Matter?
Conventional autoregressive vision models default to a raster-scan (row-major) order for flattening 2-D images into 1-D sequences. We investigate the question of whether the sequence in which image patches are presented to autoregressive vision models has an impact on their performance. We begin by outlining the datasets and models used in our empirical evaluation, followed by details of our training methodology. We then present results demonstrating that variations in patch ordering, even simple ones like column-major scans, can lead to differences in model outcomes, thereby motivating a search for more optimal, potentially learned, orderings. We define and explore six fixed patch orders: row-major, column-major, Hilbert curve, spiral, diagonal, and snake. These are visualized in Figure 1 and formally defined in Appendix F.
@@
Images captured in different contexts demonstrate varying structural biases. To study whether such datasets are susceptible to patch ordering effects to different degrees, we run experiments on two datasets: ImageNet-1K dengImageNetLargescaleHierarchical2009 (natural images) and Functional Map of the World christieFunctionalMapWorld2018 (satellite). We train on their respective training sets and report results on the validation sets.
@@
We experiment with the Vision Transformer (ViT) dosovitskiyImageWorth16x162020 , Transformer-XL (TXL) dai2019transformerxlattentivelanguagemodels , Longformer beltagy2020longformerlongdocumenttransformer , and Mamba gu2024mambalineartimesequencemodeling ; renAutoregressivePretrainingMamba2024 models. We utilized the timm implementation for the Vision Transformer (ViT) and the HuggingFace implementation for Longformer. Both were adapted with minor modifications to accommodate varying patch permutations. TXL is based on the official implementation and includes a newly introduced, learned absolute position embedding to account for changing patch orders across batches. We use ARM renAutoregressivePretrainingMamba2024 as our vision Mamba model of choice due to its training stability. All four models prepend a learnable class [CLS] token as a fixed-length representation for image classification. The [CLS] token is always retained as the first token in the sequence. All models use their respective Base configurations. Complete details about the model configurations are in Appendix C.
@@
Experiments are conducted on machines equipped with either 8$\times$80GB A100 GPUs or 4$\times$40GB A100 GPUs. All models are trained for 100 epochs the AdamW optimizer using $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay of $0.03$, and a base learning rate of $\alpha=1.0\times 10^{-4}$. Batch sizes are held constant for all runs across all model-dataset pairs (details in Appendix D). We apply cosine learning rate decay with a linear warmup over 5 epochs. For the reinforcement learning experiments introduced in Section 6, we use the same optimizer configuration but with a reduced base learning rate of $\alpha=1.0\times 10^{-5}$ and no decay.
@@ 4.1 Performance Variation Across Orderings
As anticipated, the Vision Transformer (ViT) demonstrated permutation equivariance, achieving consistent top-1 accuracy (approx. 37.5% on ImageNet-1K and 46.5% on FMoW) irrespective of patch order (Figure 2). In contrast, long-sequence models like Longformer, Mamba, and Transformer-XL (T-XL) showed performance variations dependent on patch order. T-XL and Mamba were particularly sensitive; on ImageNet-1K, T-XL’s accuracy increased by 1.92%(±0.21%) with column-major ordering and decreased by 6.43%(±0.21%) with spiral. Longformer also benefited from alternative orderings (column-major, Hilbert, snake) on ImageNet-1K, improving by up to 1.83%(±0.22%) over row-major.
Dataset characteristics altered these trends. On FMoW, Longformer’s optimal ordering shifted (e.g., diagonal increased accuracy by 1.3%(±0.22%) while column-major was detrimental). T-XL still favored column-major but also benefited from diagonal ordering on FMoW. Overall, FMoW exhibited less sensitivity to patch ordering than ImageNet, likely due to the greater homogeneity of satellite imagery compared to diverse natural images. Mamba consistently performed worse with non-row/column major orderings, likely because its fixed causal scan directions ($d\in\{\!\rightarrow,\leftarrow,\downarrow,\uparrow\}$) conflict with other patch sequences. Adapting Mamba’s scan order to the patch ordering could potentially mitigate this performance drop.
@@ 5 Learning an Optimal Patch Ordering with REOrder
In Figure 3, we observe that different patch orderings indeed lead to varying compression ratios. The row-major order, which is commonly used in vision transformers, achieves higher compression ratios, suggesting strong local redundancy. Interestingly, the Hilbert curve ordering, which aims to preserve spatial locality, shows similar compression characteristics to row-major order. In contrast, the column-major/spiral orderings exhibit lower compression ratios, indicating less local redundancy.
While this perspective offers an interesting lens through which to view patch ordering, we find that compression alone is not sufficient to predict optimal orderings for downstream tasks. For instance, while column-major ordering in Figure 3 shows lower compression ratios, it does not consistently lead to better model performance across all architectures. This motivates our subsequent exploration of learning-based approaches to discover effective patch sequences.

### v2

@@ 1 Introduction
However, this assumption does not hold in the context of modern, long-sequence models which introduce strong inductive biases such as locality beltagy2020longformerlongdocumenttransformer , recurrence dai2019transformerxlattentivelanguagemodels , or input-dependent state dynamics gu2024mambalineartimesequencemodeling that are sensitive to input ordering. Through mechanisms such as sparse attention via masking summarizing early parts of a long-sequence into latent representations, these methods are able to model long-sequences in a computationally tractable fashion. However, these design choices introduce a strong dependency on patch ordering, something that has previously been overlooked.
In this work, we demonstrate that merely swapping the row-major scan for an alternative, such as column-major or Hilbert curves, yields measurable accuracy gains across multiple long-sequence backbones. Further, we introduce REOrder, a method to learn an optimal patch ordering. REOrder first approximates which patch ordering may lead to best performance and then learns to rank patches in order of importance with reinforcement learning. Specifically, to convert the 2-D image into a 1-D sequence, we first quantify the compressibility resulting from each of six different patch orderings. Then, a Plackett-Luce ranking model is initialized with the least compressible ordering and further trained to minimize classification loss with REINFORCE. REOrder improves performance on ImageNet-1K by up to $3.01\%~(\pm 0.23\%)$ and Functional Map of the World by $13.35\%~(\pm 0.21\%)$. Code and animations are available on the project page.
@@ 2 Related Works
As established, the quadratic computational cost of full self-attention with respect to sequence length makes processing long-sequences of patches very expensive for standard Vision Transformers. To address this, significant research has focused on developing efficient transformer architectures and long-sequence models that reduce this complexity, such as Sparse Attention child2019generatinglongsequencessparse , Longformer beltagy2020longformerlongdocumenttransformer , Transformer-XL dai2019transformerxlattentivelanguagemodels , and Mamba gu2024mambalineartimesequencemodeling ; renAutoregressivePretrainingMamba2024 . While making modeling tractable, these methods often introduce specific inductive biases in how they process sequences, potentially making them sensitive to the order of the input tokens (which we discuss further in Section˜3.3). In this work, we study the effect of patch ordering on Longformer, Transformer-XL, and Mamba, demonstrating that the inductive biases inherent in these different long-sequence modeling approaches lead to substantial accuracy variance across different patch orders.
@@ 3.1 Self-attention and permutation equivariance
$\mathbf{X}=[x_{1};\dots;x_{n}]\in\mathbb{R}^{n\times d}.$
@@
$\operatorname{Attn}(\mathbf{X})=\operatorname{softmax}\!\Bigl(\tfrac{(\mathbf{X}\mathbf{W}_{q})(\mathbf{X}\mathbf{W}_{k})^{\!\top}}{\sqrt{d}}\Bigr)\,\mathbf{X}\mathbf{W}_{v},\qquad\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ (1)
where $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ are learnable query, key, and value matrices, respectively.
@@
For every permutation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$,
$\operatorname{Attn}(\mathbf{P}\mathbf{X})=\mathbf{P}\,\operatorname{Attn}(\mathbf{X}).$ (2)
Equation (2) states that self-attention is equivariant to arbitrary permutations of the patch ordering. Hence, the Vision Transformer, composed entirely of self-attention, is also permutation-equivariant. A proof for Proposition (3.1) is in Appendix˜B; a thorough analysis is conducted by Xu et. al. xuPermutationEquivarianceTransformers2024 .
@@ 3.3 Self-attention approximations and sensitivity to patch order
Recent work reduces the $\mathcal{O}(n^{2})$ cost of full attention by sparsifying the attention pattern, factoring the softmax kernel, or replacing it with a learned recurrence. These changes, however, also break the full permutation-equivariance guarantee of Eq. (2). Below we inspect three representative models. Their attention patterns are visualized in Appendix˜G.
@@
$Q=\mathbf{H}\mathbf{W}_{q},\qquad K=\tilde{\mathbf{H}}\,\mathbf{W}_{k},\qquad V=\tilde{\mathbf{H}}\,\mathbf{W}_{v}$
Let $\mathbf{P}$ permute the $L$ patches of only the current segment (but not the memory). Memory corresponds to previous segments, so its rows are fixed under permutation.
$\mathbf{P}\tilde{\mathbf{H}}=[\operatorname{SG}(\mathbf{M});\,\mathbf{P}\mathbf{H}]$
Because the second and third logit terms of $\mathbf{M}_{\text{TXL}}(i,j)$ depend on the relative index iji-j, conjugating by $\mathbf{P}$ does not commute with the soft-max:
$A^{\text{TXL}}(\mathbf{P}\mathbf{H})\neq\mathbf{P}\,A^{\text{TXL}}(\mathbf{H}).$
@@
Longformer beltagy2020longformerlongdocumenttransformer uses a sliding‑window pattern of width $w$ plus $g$ global tokens with their own projections $\mathbf{W}_{q}^{\text{(g)}},\mathbf{W}_{k}^{\text{(g)}},\mathbf{W}_{v}^{\text{(g)}}$. Because the local mask $\mathbf{M}^{\text{local}}$ fixes which $(i,j)$ pairs are valid, applying $P$ to the rows/columns re‑labels many entries as illegal $(-\infty)$ and others as legal, so
$\operatorname{softmax}(\mathbf{P}\mathbf{M}^{\text{local}}\mathbf{P}^{\!\top})\neq\mathbf{P}\,\operatorname{softmax}(\mathbf{M}^{\text{local}})\,\mathbf{P}^{\!\top}.$
The same holds for the global mask unless $\mathbf{P}$ preserves the hand‑picked global positions. Therefore, Longformer is sensitive to patch ordering by design. It converts $\mathcal{O}(n^{2})$ complexity to $\mathcal{O}(nw)$ under a rigid spatial prior.
@@
$\mathbf{h}_{t}\;=\;\mathbf{A}_{t}\,\mathbf{h}_{t-1}\;+\;\mathbf{B}_{t}\,\mathbf{x}_{t},\qquad\mathbf{y}_{t}=\mathbf{C}_{t}\,\mathbf{h}_{t},$
with inputs $\mathbf{x}_{t}$ produced by scanning the patch sequence left‑to‑right. A permutation $\mathbf{P}$ re‑orders the stream, changing $(\mathbf{B}_{t},\mathbf{C}_{t},\boldsymbol{\Delta}_{t})$ and the sequence of matrix multiplications, so the recurrence yield is different. Therefore Mamba is also permutation sensitive. Its $\mathcal{O}(n)$ complexity comes at the price of a fixed processing order.
In this work, we use ARM as introduced by Ren et. al. renAutoregressivePretrainingMamba2024 . In ARM, every Mamba layer runs four causal scans in different directions $d$ and sums their outputs before the channel‑mixing MLP.
For each $d\in\{\!\rightarrow,\leftarrow,\downarrow,\uparrow\}$ let $x_{t}^{(d)}$ be the patch encountered at time step $t$ when traversing the image in direction $d$. The direction‑specific recurrence is
$\mathbf{h}_{t}^{(d)}\;=\;\mathbf{A}_{t}^{(d)}\,\mathbf{h}_{t-1}^{(d)}+\mathbf{B}_{t}^{(d)}\,\mathbf{x}_{t}^{(d)},\qquad\mathbf{y}_{t}^{(d)}\;=\;\mathbf{C}_{t}^{(d)}\,\mathbf{h}_{t}^{(d)},$
@@
$\mathbf{y}_{t}\;=\;\sum_{d\in\{\!\rightarrow,\leftarrow,\downarrow,\uparrow\}}\mathbf{y}_{t}^{(d)}.$
Similar to Mamba, because each scan is individually directional, an arbitrary permutation $P$ of the patch order both re-orders the input streams $\{\mathbf{x}_{t}^{(d)}\}$ and alters the learned parameter sequences $\bigl\{\mathbf{A}_{t}^{(d)},\mathbf{B}_{t}^{(d)},\mathbf{C}_{t}^{(d)}\bigr\}$. Hence the mapping still violates permutation equivariance $\mathbf{y}(\mathbf{P}\mathbf{X})\neq\mathbf{P}\,\mathbf{y}(\mathbf{X})$. Further discussion in regards to the symmetry and performance implications of ARM with different orderings is discussed in  Appendix˜A
@@ 4 Does Patch Order Matter?
Conventional autoregressive vision models default to a raster-scan (row-major) order for flattening 2-D images into 1-D sequences. We investigate the question of whether the sequence in which image patches are presented to autoregressive vision models has an impact on their performance. We begin by outlining the datasets and models used in our empirical evaluation, followed by details of our training methodology. We then present results demonstrating that variations in patch ordering, even simple ones like column-major scans, can lead to differences in model outcomes, thereby motivating a search for more optimal, potentially learned, orderings. We define and explore six fixed patch orders: row-major, column-major, Hilbert curve, spiral, diagonal, and snake. These are visualized in Figure˜1 and formally defined in Appendix˜F.
@@
Images captured in different contexts demonstrate varying structural biases. To study whether such datasets are susceptible to patch ordering effects to different degrees, we run experiments on two datasets: ImageNet-1K dengImageNetLargescaleHierarchical2009 (natural images) and Functional Map of the World christieFunctionalMapWorld2018 (satellite). We train on their respective training sets and report results on the validation sets. Details on the dataset licenses are provided in  Appendix˜E
@@
We experiment with the Vision Transformer (ViT) dosovitskiyImageWorth16x162020 , Transformer-XL (TXL) dai2019transformerxlattentivelanguagemodels , Longformer beltagy2020longformerlongdocumenttransformer , and ARM model as our Mamba variant gu2024mambalineartimesequencemodeling ; renAutoregressivePretrainingMamba2024 . These three long-sequence architectures were chosen to represent distinct approaches to efficient sequence modeling: Transformer-XL employs segment-level recurrence and relative positional encodings to extend context length; Longformer reduces the quadratic attention cost through a combination of sliding-window and global attention; and Mamba introduces a structured state-space model with linear-time complexity and constant memory scaling. Together, they span the major design paradigms for long-range dependency modeling. We then test whether patch-ordering effects persist across fundamentally different inductive biases.
We utilized the timm implementation for the Vision Transformer (ViT) and the HuggingFace implementation for Longformer. Both were adapted with minor modifications to accommodate varying patch permutations. TXL is based on the official implementation and includes a newly introduced, learned absolute position embedding to account for changing patch orders across batches. We use ARM renAutoregressivePretrainingMamba2024 as our vision Mamba model of choice due to its training stability. For all of the models, the image size is $\times224$ with a patch size of $\times16$. The Transformer-XL memory length ($\mathbf{M}$) was set to 128 and the attention window size ($\mathbf{M}^{local}$) for Longformer was set to 14. All four models prepend a learnable class [CLS] token as a fixed-length representation for image classification. The [CLS] token is always retained as the first token in the sequence. All models use their respective Base configurations. Complete details about the model configurations are in Appendix˜C.
@@
Experiments are conducted on machines equipped with either $\times8$0GB A100 GPUs or $\times4$0GB A100 GPUs. We apply basic data augmentations of resizing to $\times256$, center crop to $\times224$ and then a horizontal flip with $p=0.5$. We ablate this augmentation choice in  Section˜D.3. All models are trained for 100 epochs the AdamW optimizer using $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay of $0.03$, and a base learning rate of $\alpha=1.0\times 10^{-4}$. Batch sizes are held constant for all runs across all model-dataset pairs (details in Appendix˜D). We apply cosine learning rate decay with a linear warmup over 5 epochs. For the reinforcement learning experiments introduced in Section˜6, we use the same optimizer configuration but with a reduced base learning rate of $\alpha=1.0\times 10^{-5}$ and no decay.
@@ 4.1 Performance Variation Across Orderings
As anticipated, the Vision Transformer (ViT) demonstrated permutation equivariance, achieving consistent top-1 accuracy (approx. 37.5% on ImageNet-1K and 46.5% on FMoW) irrespective of patch order (Figure 2). In contrast, long-sequence models like Longformer, Mamba, and Transformer-XL (T-XL) showed performance variations dependent on patch order. T-XL and Mamba were particularly sensitive; on ImageNet-1K, T-XL’s accuracy increased by 1.92%(±0.21 percentage points) with column-major ordering and decreased by 6.43%(±0.21 percentage points) with spiral. Longformer also benefited from alternative orderings (column-major, Hilbert, snake) on ImageNet-1K, improving by up to 1.83%(±0.22 percentage points) over row-major.
Dataset characteristics altered these trends. On FMoW, Longformer’s optimal ordering shifted (e.g., diagonal increased accuracy by 1.3%(±0.22 percentage points) while column-major was detrimental). T-XL still favored column-major but also benefited from diagonal ordering on FMoW. Overall, FMoW exhibited less sensitivity to patch ordering than ImageNet, likely due to the greater homogeneity of satellite imagery compared to diverse natural images. Mamba consistently performed worse with non-row/column major orderings, likely because its fixed causal scan directions ($d\in\{\!\rightarrow,\leftarrow,\downarrow,\uparrow\}$) conflict with other patch sequences. Adapting Mamba’s scan order to the patch ordering could potentially mitigate this performance drop.
@@ 5 Learning an Optimal Patch Ordering with REOrder
In Figure˜3, we observe that different patch orderings indeed lead to varying compression ratios. The row-major order, which is commonly used in vision transformers, achieves higher compression ratios, suggesting strong local redundancy. Interestingly, the Hilbert curve ordering, which aims to preserve spatial locality, shows similar compression characteristics to row-major order. In contrast, the column-major/spiral orderings exhibit lower compression ratios, indicating less local redundancy. This result highlights the limitations of applying a 1D compression algorithm (LZMA) to sequences derived from 2D data. While Hilbert curves preserve 2D locality better, this may not translate to higher compressibility for a 1D algorithm.
While this analysis provides a useful lens for examining patch dependencies, it produces only a weak and model-dependent correlation with downstream accuracy. That is, lower compressibility does not consistently imply improved task performance. For instance, while column-major ordering in Figure˜3 shows lower compression ratios, it does not consistently lead to better model performance across all architectures. This result is itself significant as it demonstrates that no simple, universal heuristic exists for a 1-D sequence of image tokens.
This insight motivates the second stage of our approach, in which we employ reinforcement learning to discover task-specific orderings. We use compressibility as a weak prior to give a warm start to our exploration. We assess the importance of such initializations in  Section˜D.2.